### v2

# Quantum automorphism groups of 2-graphs

Abstract.
We formulate a notion of the quantum automorphism groups of $2$-graphs. We show that two isomorphic $2$-graphs have isomorphic quantum automorphism groups making the notion well-defined. We compute the quantum automorphism groups of a few $2$-graphs. It is observed that quantum automorphism group of a $1$-graph can be realized as the quantum automorphism group of a $2$-graph.
Key words and phrases:
higher rank graphs, quantum automorphism groups, quantum groups, $C^{*}$-algebra
1991 Mathematics Subject Classification:
46L67
1. Introduction
Higher rank graphs (synonymously rank $k$ graphs or $k$-graphs) are natural generalization of graphs. Roughly speaking, a $k$-graph $\Lambda$ is a countable small category equipped with a “degree map” $d:\Lambda\rightarrow\mathbb{N}^{k}$ with the morphisms admitting a unique factorization property. Since its inception, it has played a key role in the realm of noncommutative geometry and topology. In particular, it helps in the proper generalization of Cuntz-Krieger $C^{\ast}$-algebras to higher-rank graph $C^{\ast}$-algebras. There is a formidable literature on higher rank graphs and their $C^{\ast}$-algebras (see [NYJM], [Steger1], [Steger2], [Sims], [Sims1], [Sims2], [Hazzlewood1] to name a few). Pioneered by S. Wang ([Wang]), the quantum symmetry of various finite dimensional or infinite dimensional classical or quantum objects has been quite extensively studied by many authors in recent years (e.g. [Goswami], [Banica], [Hajac], [Skalski], [Asfaq], [Voigt2] [Vaes]). Such quantum symmetry is now well established as the generalized symmetry in the world of noncommutative geometry or topology. Among the finite structures, the quantum automorphism group of finite graphs (see [Schmidt], [Fulton]), at least for small number of vertices, are reasonably well understood by now. Moreover, the notion of the quantum automorphism groups of graphs has found its way into quantum information theory (see [Qinfo]) as well as some concrete physical models involving graphs (see [Potts]). Therefore, higher rank graphs being a generalization of usual graphs, it is quite natural to consider the notion of quantum automorphism groups of higher rank graphs. Recall that the quantum automorphism groups of $C^{\ast}$-algebras are typically defined to be the universal objects of certain category of compact quantum groups acting faithfully on the $C^{\ast}$-algebras. Existence of the universal objects in such categories are often tricky and requires careful consideration. In this paper, we define the notion of “the” quantum automorphism group of rank $2$ graphs. We restrict our attention to the class of $2$-graphs as they have an equivalent formulation in terms of a simpler combinatorial data circumventing the category theoretic machinery. Briefly speaking, a $2$-graph can be described by a triple consisting of two $1$-graphs on the same vertex set having commuting vertex matrices and a bijection between certain sets. We would like to mention that we restrict our attention to the case where the common vertex set has finitely many points and the individual graphs are without multiple edge, source or sink to be able to work in the category of compact quantum groups. Then the quantum automorphism group of such a triple is naturally defined. But one has to be careful as the isomorphism class of $2$-graphs depend on a natural notion of equivalence between triples. We consider the right notion of two equivalent triples and prove that the quantum automorphism groups of such equivalent triples are isomorphic. This allows us to call the quantum automorphism group of a triple “the” quantum automorphism group of the corresponding $2$-graph. We then compute the quantum automorphism group of some $2$-graphs. It is clear from the definition that in order to find the quantum automorphism group of a $2$-graph to be genuine one has to at least start with a triple which has two constituent graphs admitting genuine quantum automorphisms. But that is naturally not sufficient though as seen in Section 4 Example (b). A particular class of $2$-graphs are of importance where the constituent graphs of the triple are two copies of the same graph. In such cases if the bijection is the trivial one then obviously one gets the quantum automorphism group of the $1$-graph as the quantum automorphism group of the corresponding $2$-graph. In that way one can easily ensure genuine quantum symmetry by choosing the $1$-graph with genuine quantum symmetry. However, the case of non-trivial bijection is clearly more interesting. We have an example with a non-trivial bijection where the quantum automorphism group of the resulting $2$-graph is still genuine. Also we exhibit one example where we have two copies of the complete graph on $4$-vertices as the constituent graphs of the triple, but there is a choice of a bijection so that the quantum automorphism group of the corresponding triple is not only classical but also it is trivial. We end this paper with a brief discussion on possible problems to explore in future. We have also added an Appendix section where a brief sketch of the proof of the fact that the isomorphism classes of $2$-graphs depend on an equivalence class of triples has been included.
2. Preliminaries
2.1. Higher rank graphs
Let us recall the definition of a higher rank graph briskly. We shall be concerned about the case when $k=2$. But for the sake of completeness we define a $k$-graph anyway. For details, the reader is referred to [NYJM].
Definition 2.2.
(see Definition $1.1$ of [NYJM]) A $k$-graph (rank $k$ graph or higher rank graph) $(\Lambda,d)$ consists of a countable small category $\Lambda$ (with range and source maps $r$ and $s$ respectively) together with a functor $d:\Lambda\rightarrow\mathbb{N}^{k}$ satisfying the factorisation property: for every morphism $\lambda\in\Lambda$ and $m,n\in\mathbb{N}^{k}$ with $d(\lambda)=m+n,$ there are unique elements $\mu,\nu\in\Lambda$ such that $\lambda=\mu\nu$ and $d(\mu)=m,\ d(\nu)=n.$
Remark 2.3.
For $n\in\mathbb{N}^{k}$, one denotes the set of morphisms of degree $n$ (i.e. $d^{-1}(n)$) by $\Lambda^{n}$. The unique factorization property enables one to identify $\Lambda^{0}$ with ${\rm obj}(\Lambda)$. We shall restrict ourselves to the case when $\Lambda^{0}$ is finite.
Definition 2.4.
(see Definition $1.9$ of [NYJM]) Let $f:\mathbb{N}^{l}\rightarrow\mathbb{N}^{k}$ be a monoid morphism. If $(\Lambda,d)$ is a $k$-graph we may form the $l$-graph $f^{\ast}(\Lambda)$ as follows: $f^{\ast}(\Lambda)=\{(\lambda,n):d(\lambda)=f(n)\}\ {\rm with}\ d(\lambda,n)=n,s(\lambda,n)=s(\lambda)\ {\rm and}\ r(\lambda,n)=r(\lambda)$.
Now let us specialize to the case where $k=2$. We shall take an alternative equivalent view point of $2$-graphs which will be amenable to the study of quantum symmetry. This view point is described in Section $6$ of [NYJM]. Here we shall give the bare bones. Roughly speaking, any $2$-graph can be constructed from two $1$-graphs with common set of vertices and commuting vertex matrices. Again we remind the reader that in our case, since ${\rm obj}(\Lambda)$ is finite, the number of common vertices are finite.
$2$-graph: Let $\mathcal{G}_{1}=(V,E_{1})$, $\mathcal{G}_{2}=(V,E_{2})$ be two $1$-graphs with common set of vertices $V$ such that the corresponding vertex matrices and commute. The commutation of $A_{1}$ and $A_{2}$ guarantees the existence of a bijection $\theta:E_{1}\star E_{2}\rightarrow E_{2}\star E_{1}$ where
$E_{1}\star E_{2}=\{(e^{1}_{\alpha},e^{2}_{\beta})\in E_{1}\times E_{2}|s(e^{1}_{\alpha})=t(e^{2}_{\beta})\},$
$E_{2}\star E_{1}=\{(e^{2}_{\alpha},e^{1}_{\beta})\in E_{2}\times E_{1}|s(e^{2}_{\alpha})=t(e^{1}_{\beta})\}.$
We call the data $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ a triple. Given such a triple, one can construct a $2$-graph $\Lambda$ in the following way:
The set ${\rm obj}(\Lambda)$ or $\Lambda^{(0,0)}$ is the set of common vertices of the graphs $\mathcal{G}_{1},\mathcal{G}_{2}$. The morphisms of degree $(1,0)$ and $(0,1)$ are determined by the edge sets $E_{1},E_{2}$. The higher degree morphisms are built step by step using the bijection $\theta$. For details of the construction, we again refer the readers to the Section 6 of [NYJM]. Conversely given a $2$-graph $\Lambda$, one can construct a triple. Roughly one takes the set $\Lambda^{(0,0)}$ to be the set of common vertices $V$. Then the morphisms $\Lambda^{(1,0)}$ and $\Lambda^{(0,1)}$ together with the range and source maps defines the constituent $1$-graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. The bijection $\theta$ is then built out of the morphisms $\Lambda^{(1,1)}$ using the unique factorization property of the morphisms. A triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ corresponding to a $2$-graph $\Lambda$ will be called the defining triple of $\Lambda$.
Remark 2.5.
(1) A bijection between the sets $E_{1}\star E_{2}$ and $E_{2}\star E_{1}$ is guaranteed precisely because the vertex matrices commute. Also, $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$Rightarrow
$s(e^{1}_{\alpha})=t(e^{2}_{\beta});\ s(e^{2}_{\mu})=t(e^{1}_{\nu});$
$t(e^{1}_{\alpha})=t(e^{2}_{\mu});\ s(e^{2}_{\beta})=s(e^{1}_{\nu}).$
We shall refer to a $2$-graph by the triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$. Observe that given the bijection $\theta$, one has a natural vector space isomorphism $\theta_{\ast}:C(E_{1}\star E_{2})\rightarrow C(E_{2}\star E_{1})$ given by $\theta_{\ast}(f)(e^{2}_{\mu},e^{1}_{\nu}):=f(\theta^{-1}(e^{2}_{\mu},e^{1}_{\nu}))$ for $f\in C(E_{1}\star E_{2})$.
(2) We also mention that we restrict our attention to the $2$-graphs where the individual $1$-graphs are without multiple edges, sinks or sources for technical reasons.
Remark 2.6.
A case of particular interest is when we have two copies of a $1$-graph say $\mathcal{G}=(V,E)$ and the identity bijection $\theta$ from $E\star E$ to $E\star E$. In that case the resulting $2$-graph is isomorphic to the $2$-graph $f^{\ast}(\mathcal{G})$ where $f:\mathbb{N}^{2}\rightarrow\mathbb{N}$ is the monoid morphism $f(m,n)=m+n$.
Now let $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ be two triples. We denote the common vertex set of the triples by $V_{1},V_{2}$ respectively; the vertex matrices of $\mathcal{G}_{i}$ by $A_{i}$ and vertex matrices of $\mathcal{H}_{i}$ by $B_{i}$ for $i=1,2$ respectively; the edge sets of $\mathcal{G}_{i}$ and $\mathcal{H}_{i}$ by $E_{i}$ and $F_{i}$ for $i=1,2$ respectively.
Definition 2.7.
Two triples $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ as above are said to be equivalent if the following happens:
(i) There is a bijection $T:V_{1}\rightarrow V_{2}$. This means effectively we can assume that all the $1$-graphs involved share a common set of vertices $V$ and $T$ is a permutation of the vertices so that we have an invertible linear map $T_{\ast}:C(V)\rightarrow C(V)$.
(ii) If we denote the permutation matrix of the linear map $T_{\ast}$ again by $T$ with respect to standard bases of $C(V)$ then $TA_{2}=A_{1}T$ and $TB_{2}=B_{1}T$.
(iii) Consider the map $T^{\otimes 4}\equiv T_{\ast}\otimes T_{\ast}\otimes T_{\ast}\otimes T_{\ast}:C(V\times V\times V\times V)\rightarrow C(V\times V\times V\times V)$. Thanks to condition (ii), $T^{\otimes 4}$ is also a bijection between both $C(E_{1}\star F_{1})$, $C(E_{2}\star F_{2})$ and $C(F_{1}\star E_{1})$, $C(F_{2}\star E_{2})$. We have the following commutative diagram:
{tikzcd}
Remark 2.8.
We abbreviate the last condition of the above definition as $\theta$-compatibility. The isomorphism class of $2$-graphs depend only on the equivalence class of triples as defined above. We have included a brief sketch of the proof. We believe that it must be known. But since we could not find an explicit mention of this result in the literature, we have decided to keep the result in the Appendix (see Theorem A.1).
2.9. Compact quantum groups and their actions
The term quantum group is quite general. We make it clear that we work with Compact quantum groups (CQG in short) ala S.L. Woronowicz ([Woro]). The literature on CQG and their actions on $C^{\ast}$-algebras are aplenty by now. The reader can refer to the articles [DeCom], [Soltan] for details on the actions of CQG’s on $C^{\ast}$-algebras. Here we briefly recall what is relevant for us. The following are mainly taken from [edin].
Definition 2.10.
A compact quantum group (CQG) $\mathbb{G}$ is a pair $(C(\mathbb{G}),\Delta_{\mathbb{G}})$ such that $C(\mathbb{G})$ is a unital $C^{\ast}$-algebra and $\Delta_{\mathbb{G}}:C(\mathbb{G})\rightarrow C(\mathbb{G})\otimes C(\mathbb{G})$ is a unital $C^{\ast}$- homomorphism satisfying
(i) $({\rm id}\otimes\Delta_{\mathbb{G}})\circ\Delta_{\mathbb{G}}=(\Delta_{\mathbb{G}}\otimes{\rm id})\circ\Delta_{\mathbb{G}}.$
(ii) ${\rm Span}\{\Delta_{\mathbb{G}}(C(\mathbb{G}))(1\otimes C(\mathbb{G}))\}$ and ${\rm Span}\{\Delta_{\mathbb{G}}(C(\mathbb{G}))(C(\mathbb{G})\otimes 1)\}$ are dense in $C(\mathbb{G})\otimes C(\mathbb{G})$.
Remark 2.11.
Given a CQG $\mathbb{G}$, there is a canonical dense Hopf $\ast$-algebra $C(\mathbb{G})_{0}$ in $C(\mathbb{G})$ on which an antipode $\kappa$ and counit $\epsilon$ are defined. Given two CQG’s $\mathbb{G}_{1}$ and $\mathbb{G}_{2}$, a CQG morphism between them is a unital $C^{\ast}$-homomorphism $\pi:C(\mathbb{G}_{1})\rightarrow C(\mathbb{G}_{2})$ such that $(\pi\otimes\pi)\circ\Delta_{\mathbb{G}_{1}}=\Delta_{\mathbb{G}_{2}}\circ\pi.$
Definition 2.12.
Given a (unital) $C^{\ast}$-algebra ${\cal C}$, a CQG $(C(\mathbb{G}),\Delta_{\mathbb{G}})$ is said to act faithfully on ${\cal C}$ if there is a unital $C^{\ast}$-homomorphism $\alpha:{\cal C}\rightarrow{\cal C}\otimes C(\mathbb{G})$ satisfying
(i) $(\alpha\otimes{\rm id})\circ\alpha=({\rm id}\otimes\Delta_{\mathbb{G}})\circ\alpha$.
(ii) Span$\{\alpha({\cal C})(1\otimes\Delta_{\mathbb{G}})\}$ is dense in ${\cal C}\otimes C(\mathbb{G})$.
(iii) The $\ast$-algebra generated by the set $\{(\omega\otimes{\rm id})\circ\alpha({\cal C}):\omega\in{\cal C}^{\ast}\}$ is norm-dense in $C(\mathbb{G})$.
Definition 2.13.
(Def 2.1 of [Bichon]) Given a unital $C^{\ast}$-algebra ${\cal C}$, quantum automorphism group of ${\cal C}$ is a CQG $(C(\mathbb{G}),\Delta_{\mathbb{G}})$ acting faithfully on ${\cal C}$ satisfying the following universal property:
If $\mathbb{B}$ is any CQG acting faithfully on ${\cal C}$ with the action $\beta$, there is a surjective CQG morphism $\pi:C(\mathbb{G})\rightarrow C(\mathbb{B})$ such that $({\rm id}\otimes\pi)\circ\alpha=\beta$.
Remark 2.14.
In general the universal object might fail to exist in the above category. To ensure the existence one generally assumes that the action preserves some fixed state on the $C^{\ast}$-algebra. However for certain finite dimensional $C^{\ast}$-algebras such a state preserving condition is automatic. We would not go into further details as we are not going to use it in this paper. For further details, the reader might consult [Wang].
Example 2.15.
If we take the space of $n$ points $X_{n}$ then the quantum automorphism group of the $C^{\ast}$-algebra $C(X_{n})$ is given by the CQG (denoted by $S_{n}^{+}$) whose underlying $C^{\ast}$-algebra is the universal $C^{\ast}$ algebra generated by $\{q_{ij}\}_{i,j=1,\ldots,n}$ satisfying the following relations (see Theorem 3.1 of [Wang]):
$q_{ij}^{2}=q_{ij},\quad q_{ij}^{\ast}=q_{ij},\quad\sum_{k=1}^{n}q_{ik}=\sum_{k=1}^{n}q_{ki}=1,\quad\text{ for }i,j=1,\ldots,n.$
Any $n\times n$ matrix $((q_{ij}))_{i,j=1,\ldots,n}$ with $q_{ij}$’s in a $C^{\ast}$-algebra satisfying the above relations is called a magic unitary. The coproduct on the generators is given by $\Delta(q_{ij})=\sum_{k=1}^{n}q_{ik}\otimes q_{kj}$. It is worth mentioning that the CQG $S_{n}^{+}$ is a particular type of CQG known as the compact matrix pseudogroup (see [Woro]). The $\ast$-algebra generated by $\{q_{ij}\}_{i,j=1,\ldots,n}$ is the canonical dense Hopf $\ast$-algebra and the antipode is given on the generators by $\kappa(q_{ij})=q_{ji}$.
2.16. Quantum automorphism group of finite graphs
Recall the definition of finite, directed graph $\mathcal{G}=(V=\{v_{1},\ldots,v_{n}\},E=\{e_{1},\ldots,e_{m}\})$ without multiple edge and the CQG $S_{n}^{+}$. We shall call a finite, directed graph a $1$-graph to be consistent with higher rank graph terminology from now on.
Definition 2.17.
The quantum automorphism group of a graph $\mathcal{G}$ without multiple edges (to be denoted by ${\rm Aut}^{+}(\mathcal{G})$) is defined to be the quotient $S^{+}_{n}/(AD-DA)$, where $A=((q_{ij}))_{i,j=1,\ldots,n}$, and $D$ is the vertex matrix of $\mathcal{G}$. The coproduct on the generators is given by $\Delta(q_{ij})=\sum_{k=1}^{n}q_{ik}\otimes q_{kj}$.
For the classical automophism group ${\rm Aut}(\mathcal{G})$, the commutative CQG $C({\rm Aut}(\mathcal{G}))$ is generated by $q_{ij}$’s where $q_{ij}$ is a function on $S_{n}$ taking value $1$ on the permutation which sends $i$-th vertex to $j$-th vertex and takes the value zero on other elements of the group. It is a quantum subgroup of ${\rm Aut}^{+}(\mathcal{G})$. The surjective CQG morphism $\pi:C({\rm Aut}^{+}(\mathcal{G}))\rightarrow C({\rm Aut}(\mathcal{G}))$ sends the generators to generators.
Theorem 2.18.
(Lemma 3.1.1 of [Fulton]) The underlying $C^{\ast}$-algebra of the quantum automorphism group ${\rm Aut}^{+}(\mathcal{G})$ of a finite graph $\mathcal{G}$ with $m$ edges and $n$ vertices (without multiple edge) is the universal $C^{\ast}$-algebra generated by $\{q_{ij}\}_{i,j=1,\ldots,n}$ satisfying the following relations:
  1. (1)
    $q_{ij}^{\ast}=q_{ij},\quad q_{ij}q_{ik}=\delta_{jk}q_{ij},\quad q_{ji}q_{ki}=\delta_{jk}q_{ji},\text{ for }1\leq i,j,k\leq n;$
  2. (2)
    $\sum_{l=1}^{n}q_{il}=\sum_{l=1}^{n}q_{li}=1,\text{ for }1\leq i\leq n;$
  3. (3)
    $q_{s(e_{j})i}q_{t(e_{j})k}=q_{t(e_{j})k}q_{s(e_{j})i}=q_{is(e_{j})}q_{kt(e_{j})}=q_{kt(e_{j})}q_{is(e_{j})}=0,$ whenever $e_{j}\in E$ and $(i,k)\not\in E.$
The coproduct on the generators is given by $\Delta(q_{ij})=\sum_{k=1}^{n}q_{ik}\otimes q_{kj}$. The $C^{\ast}$-action on the graph is given by $\alpha(\delta_{i})=\sum_{j}\delta_{j}\otimes q_{ji}$, where $\delta_{i}$ is the function which takes value $1$ on $i$-th vertex and zero elsewhere.
3. Main Results
To define the quantum automorphism group of a $2$-graph, we shall define the notion of the quantum automorphism group of a triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ first. Then we shall show that the quantum automorphism groups of two such equivalent triples are isomorphic. Then as the isomorphism class of a $2$-graph depends on the equivalence class of its defining triple, one can define the quantum automorphism group of a $2$-graph to be the quantum automorphism group of its defining triple unambiguously. To that end let $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ be a triple where we remind the readers that $\mathcal{G}_{i}=(V,E_{i})$ are $1$-graphs (without any multiple edge, source or sink) with common set of finitely many vertices $V$ and edge sets $E_{i}$’s for $i=1,2$ respectively; Let the vertex matrices of $\mathcal{G}_{i}$ be $A_{i}$ for $i=1,2$ respectively such that $A_{1}A_{2}=A_{2}A_{1}$. Given a CQG $\mathbb{G}$ having an action $\alpha:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ given by $\alpha(\delta_{v_{i}})=\sum_{j}\delta_{v_{j}}\otimes q_{ji}$ for $q_{ji}\in C(\mathbb{G})$ such that $((q_{ij}))$ commutes with $A_{1}$ and $A_{2}$, there are corepresentations $\beta_{i}$ on $C(E_{i})$ given by
$\beta_{i}(\delta_{e^{i}_{\alpha}})=\sum_{\beta}\delta_{e^{i}_{\beta}}\otimes q_{t(e^{i}_{\beta})t(e^{i}_{\alpha})}q_{s(e^{i}_{\beta})s(e^{i}_{\alpha})}.$
Then one can consider tensor product corepresentations $\beta_{1}\otimes\beta_{2}$, $\beta_{2}\otimes\beta_{1}$ on $C(E_{1}\times E_{2})$ and $C(E_{2}\times E_{1})$ respectively. Now note that $C(E_{1}\star E_{2})$ is a linear subspace of $C(E_{1}\times E_{2})$ with a basis for $C(E_{1}\star E_{2})$ given by $\{\delta_{e^{1}_{\alpha}}\otimes\delta_{e^{2}_{\beta}}:s(e^{1}_{\alpha})=t(e^{2}_{\beta})\}$. Similarly $C(E_{2}\star E_{1})$ is a linear subspace of $C(E_{2}\times E_{1})$.
Lemma 3.1.
The corepresentation $\beta_{1}\otimes\beta_{2}$ leaves the subspace $C(E_{1}\star E_{2})$ invariant.
Proof.
By definition of $\beta_{1}\otimes\beta_{2}$, we have the following:
$(\beta_{1}\otimes\beta_{2})(\delta_{e^{1}_{\alpha}}\otimes\delta_{e^{2}_{\beta}})=\sum\delta_{e^{1}_{\mu}}\otimes\delta_{e^{2}_{\nu}}\otimes q_{t(e^{1}_{\mu})t(e^{1}_{\alpha})}q_{s(e^{1}_{\mu})s(e^{1}_{\alpha})}q_{t(e^{2}_{\nu})t(e^{2}_{\beta})}q_{s(e^{2}_{\nu})s(e^{2}_{\beta})}.$
Now suppose $s(e^{1}_{\alpha})=t(e^{2}_{\beta})$ so that $\delta_{e^{1}_{\alpha}}\otimes\delta_{e^{2}_{\beta}}\in C(E_{1}\star E_{2})$. Then for $(e^{1}_{\mu},e^{2}_{\nu})$ such that $s(e^{1}_{\mu})\neq t(e^{2}_{\nu})$, we have $q_{t(e^{1}_{\mu})t(e^{1}_{\alpha})}q_{s(e^{1}_{\mu})s(e^{1}_{\alpha})}q_{t(e^{2}_{\nu})t(e^{2}_{\beta})}q_{s(e^{2}_{\nu})s(e^{2}_{\beta})}=0$ which proves the lemma. $\Box$
Remark 3.2.
(i) By the same reasoning one can show that $\beta_{2}\otimes\beta_{1}$ leaves $C(E_{2}\star E_{1})$ invariant. We denote the restriction of $\beta_{1}\otimes\beta_{2}$ on $C(E_{1}\star E_{2})$ by $\beta$ and the restriction of $\beta_{2}\otimes\beta_{1}$ on $C(E_{2}\star E_{1})$ by $\beta^{{\rm op}}$.
(ii) The corepresentations $\beta$, $\beta^{\rm op}$ are nothing but the restrictions of the tensor product corepresentation $\alpha^{\otimes 4}:=\alpha\otimes\alpha\otimes\alpha\otimes\alpha$ of $\mathbb{G}$ on $C(V\times V\times V\times V)$ to the subspaces $C(E_{1}\star E_{2})$ and $C(E_{2}\star E_{1})$ respectively.
Now recall the vector space isomorphism $\theta_{\ast}$ corresponding to the bijection $\theta$.
Definition 3.3.
A CQG $\mathbb{G}$ is said to act on a triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$, if there is a faithful action $\alpha:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ given by $\alpha(\delta_{v_{i}})=\sum_{j}\delta_{v_{j}}\otimes q_{ji}$ such that
(i) The matrix $((q_{ij}))$ commutes with the vertex matrices of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ so that we have the corresponding linear maps $\beta$ and $\beta^{{\rm op}}$ on $C(E_{1}\star E_{2})$ and $C(E_{2}\star E_{1})$ respectively.
(ii) The following diagram commutes:
{tikzcd}
Generators and relations: Let $\mathbb{G}$ be a CQG acting on a triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ so that (assume there are $n$-vertices) $\alpha:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ is a faithful action given by $\alpha(\delta_{v_{i}})=\sum\delta_{v_{j}}\otimes q_{ji}$. Then $C(\mathbb{G})$ is generated by $\{q_{ij}\}_{i,j=1,\ldots,n}$. Now $q_{ij}$’s satisfy all the relations of ${\rm Aut}^{+}(\mathcal{G}_{i})$ for $i=1,2$ as described in the Theorem 2.18. Let us figure out the relations imposed by the commutative diagram in the Definition (3.3). To that end we start with the easy observation that if $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$, then $\theta_{\ast}(\delta_{e^{1}_{\alpha}}\otimes\delta_{e^{2}_{\beta}})=\delta_{e^{2}_{\mu}}\otimes\delta_{e^{1}_{\nu}}$. Let $(e^{1}_{\alpha},e^{2}_{\beta})\in E_{1}\star E_{2}$ such that $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$ for some $e^{2}_{\mu},e^{1}_{\nu}\in E_{2}\star E_{1}$. Then
$\beta^{\rm op}\circ\theta_{\ast}(\delta_{e^{1}_{\alpha}}\otimes\delta_{e^{2}_{\beta}})$
$=\beta^{\rm op}(\delta_{e^{2}_{\mu}}\otimes\delta_{e^{1}_{\nu}})$
$=\sum_{(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})\in E_{2}\star E_{1}}\delta_{e^{2}_{\mu^{\prime}}}\otimes\delta_{e^{1}_{\nu^{\prime}}}\otimes q_{t(e^{2}_{\mu^{\prime}})t(e^{2}_{\mu})}q_{s(e^{2}_{\mu^{\prime}})s(e^{2}_{\mu})}q_{t(e^{1}_{\nu^{\prime}})t(e^{1}_{\nu})}q_{s(e^{1}_{\nu^{\prime}})s(e^{1}_{\nu})}$
(3.1) $=\sum_{(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})\in E_{2}\star E_{1}}\delta_{e^{2}_{\mu^{\prime}}}\otimes\delta_{e^{1}_{\nu^{\prime}}}\otimes q_{t(e^{2}_{\mu^{\prime}})t(e^{2}_{\mu})}q_{s(e^{2}_{\mu^{\prime}})s(e^{2}_{\mu})}q_{s(e^{1}_{\nu^{\prime}})s(e^{1}_{\nu})}.$
On the other hand
$(\theta_{\ast}\otimes{\rm id})\circ\beta(\delta_{e^{1}_{\alpha}}\otimes\delta_{e^{2}_{\beta}})$
$=(\theta_{\ast}\otimes{\rm id})\left(\sum_{(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})\in E_{1}\star E_{2}}\delta_{e^{1}_{\alpha^{\prime}}}\otimes\delta_{e^{2}_{\beta^{\prime}}}\otimes q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{t(e^{2}_{\beta^{\prime}})t(e^{2}_{\beta})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}\right)$
Now let $\theta(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})=(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$ and observe that as $\theta$ is a bijection, if $(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})$ vary over all $E_{1}\star E_{2}$, $(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$ vary over all of $E_{2}\star E_{1}$ so that
$(\theta_{\ast}\otimes{\rm id})\circ\beta(\delta_{e^{1}_{\alpha}}\otimes\delta_{e^{2}_{\beta}})$
(3.2) $=\sum_{(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})\in E_{2}\star E_{1}}\delta_{e^{2}_{\mu^{\prime}}}\otimes\delta_{e^{1}_{\nu^{\prime}}}\otimes q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}$
Comparing coefficients, we get for all $(e^{1}_{\alpha},e^{2}_{\beta})\in E_{1}\star E_{2}$ and all $(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})\in E_{2}\star E_{1}$ with $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$ and $\theta(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})=(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$,
(3.3) $q_{t(e^{2}_{\mu^{\prime}})t(e^{2}_{\mu})}q_{s(e^{2}_{\mu^{\prime}})s(e^{2}_{\mu})}q_{s(e^{1}_{\nu^{\prime}})s(e^{1}_{\nu})}=q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}.$
Observation: Recall that $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$ and $\theta(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})=(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$Rightarrow $t(e^{2}_{\mu^{\prime}})=t(e^{1}_{\alpha^{\prime}})$, $t(e^{2}_{\mu})=t(e^{1}_{\alpha})$, $s(e^{1}_{\nu^{\prime}})=s(e^{2}_{\beta^{\prime}})$ and $s(e^{1}_{\nu})=s(e^{2}_{\beta})$. Therefore the relations (3.3) have a few equivalent variations as $q_{t(e^{2}_{\mu^{\prime}})t(e^{2}_{\mu})}=q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}$ and $q_{s(e^{1}_{\nu^{\prime}})s(e^{1}_{\nu})}=q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}$.
Definition 3.4.
The quantum automorphism group of a triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ (to be denoted by ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$) is defined to be the universal CQG in the category of CGQ’s acting on the triple in the sense of Definition 3.3.
Theorem 3.5.
Let $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ be a triple such that $\mathcal{G}_{1},\mathcal{G}_{2}$ have a common set of vertices $V$ with $n$-points and both the $1$-graphs do not admit any multiple edge, source or sinks. Then ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ exists. The underlying $C^{\ast}$-algebra of ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ is the universal $C^{\ast}$-algebra generated by $n^{2}$-elements $q_{ij}$’s such that
(i) The elements $q_{ij}$’s satisfy the relations of ${\rm Aut}^{+}(\mathcal{G}_{i})$ for $i=1,2$.
(ii) The elements $q_{ij}$’s satisfy the relations (3.3).
The coproduct is given on the generators by $\Delta(q_{ij})=\sum_{k=1}^{n}q_{ik}\otimes q_{kj}$. If the usual basis for $C(V)$ is given by $\{\delta_{i}\}_{i=1}^{n}$, the coaction on $C(V)$ is given by $\alpha(\delta_{i})=\sum_{j=1}^{n}\delta_{j}\otimes q_{ji}$ for $i=1,\ldots,n$.
Proof.
Note that it suffices to show that the ideal generated by the relations (3.3) is a Woronowicz ideal i.e. for all $(e^{1}_{\alpha},e^{2}_{\beta})\in E_{1}\star E_{2}$ and all $(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})\in E_{2}\star E_{1}$ with $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$ and $\theta(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})=(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$,
$\Delta\Big{(}q_{t(e^{2}_{\mu^{\prime}})t(e^{2}_{\mu})}q_{s(e^{2}_{\mu^{\prime}})s(e^{2}_{\mu})}q_{s(e^{1}_{\nu^{\prime}})s(e^{1}_{\nu})}\Big{)}=\Delta\Big{(}q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}\Big{)}.$
The left hand side is equal to
$\sum_{i,j,k\in V}q_{t(e^{2}_{\mu^{\prime}})i}q_{s(e^{2}_{\mu^{\prime}})j}q_{s(e^{1}_{\nu^{\prime}})k}\otimes q_{it(e^{2}_{\mu})}q_{js(e^{2}_{\mu})}q_{ks(e^{1}_{\nu})}.$
Now for $(j,i)\notin E_{2}$, $q_{t(e^{2}_{\mu^{\prime}})i}q_{s(e^{2}_{\mu^{\prime}})j}=0$. As none of the graphs $\mathcal{G}_{i}$ admits multiple edges, the last summation reduces to
$\sum_{f^{2}_{\mu^{\prime\prime}}\in E_{2},k\in V}q_{t(e^{2}_{\mu^{\prime}})t(f^{2}_{\mu^{\prime\prime}})}q_{s(e^{2}_{\mu^{\prime}})s(f^{2}_{\mu^{\prime\prime}})}q_{s(e^{1}_{\nu^{\prime}})k}\otimes q_{t(f^{2}_{\mu^{\prime\prime}})t(e^{2}_{\mu})}q_{s(f^{2}_{\mu^{\prime\prime}})s(e^{2}_{\mu})}q_{ks(e^{1}_{\nu})}.$
Since $\mathcal{G}_{1}$ is without source or sink, there is an edge $f^{1}_{\alpha^{\prime\prime}}\in E_{1}$ such that $s(f^{2}_{\mu^{\prime\prime}})=t(f^{1}_{\alpha^{\prime\prime}})$. In particular this implies that $(f^{2}_{\mu^{\prime\prime}},f^{1}_{\alpha^{\prime\prime}})\in E_{2}\star E_{1}$. Replacing $s(e^{2}_{\mu^{\prime}})$ by $t(e^{1}_{\nu^{\prime}})$ and using the fact that $q_{t(e^{1}_{\nu^{\prime}})t(f^{1}_{\alpha^{\prime\prime}})}q_{s(e^{1}_{\nu^{\prime}})k}=0$ for $k\neq s(f^{1}_{\alpha^{\prime\prime}})$, the last summation reduces to
$\qquad\sum_{\mathclap{(f^{2}_{\mu^{\prime\prime}},f^{1}_{\alpha^{\prime\prime}})\in E_{2}\star E_{1}}}q_{t(e^{2}_{\mu^{\prime}})t(f^{2}_{\mu^{\prime\prime}})}q_{t(e^{1}_{\nu^{\prime}})t(f^{1}_{\alpha^{\prime\prime}})}q_{s(e^{1}_{\nu^{\prime}})s(f^{1}_{\alpha^{\prime\prime}})}\otimes q_{t(f^{2}_{\mu^{\prime\prime}})t(e^{2}_{\mu})}q_{t(f^{1}_{\alpha^{\prime\prime}})s(e^{2}_{\mu})}q_{s(f^{1}_{\alpha^{\prime\prime}})s(e^{1}_{\nu})}$
which is equal to
$\qquad\sum_{\mathclap{(f^{2}_{\mu^{\prime\prime}},f^{1}_{\alpha^{\prime\prime}})\in E_{2}\star E_{1}}}q_{t(e^{2}_{\mu^{\prime}})t(f^{2}_{\mu^{\prime\prime}})}q_{s(e^{2}_{\mu^{\prime}})s(f^{2}_{\mu^{\prime\prime}})}q_{s(e^{1}_{\nu^{\prime}})s(f^{1}_{\alpha^{\prime\prime}})}\otimes q_{t(f^{2}_{\mu^{\prime\prime}})t(e^{2}_{\mu})}q_{s(f^{2}_{\mu^{\prime\prime}})s(e^{2}_{\mu})}q_{s(f^{1}_{\alpha^{\prime\prime}})s(e^{1}_{\nu})}.$
Let $(f^{1}_{\gamma^{\prime\prime}},f^{2}_{\delta^{\prime\prime}})\in E_{1}\star E_{2}$ be such that $\theta(f^{1}_{\gamma^{\prime\prime}},f^{2}_{\delta^{\prime\prime}})=(f^{2}_{\mu^{\prime\prime}},f^{1}_{\alpha^{\prime\prime}})$. Then $\theta$ being a bijection, $(f^{1}_{\gamma^{\prime\prime}},f^{2}_{\delta^{\prime\prime}})$ vary over all of $E_{1}\star E_{2}$ as ($f^{2}_{\mu^{\prime\prime}},f^{1}_{\alpha^{\prime\prime}})$ vary over all of $E_{2}\star E_{1}$. Therefore using the relations (3.3), the last summation becomes
$\quad\qquad\sum_{\mathclap{(f^{1}_{\gamma^{\prime\prime}},f^{2}_{\delta^{\prime\prime}})\in E_{1}\star E_{2}}}q_{t(e^{1}_{\alpha^{\prime}})t(f^{1}_{\gamma^{\prime\prime}})}q_{s(e^{1}_{\alpha^{\prime}})s(f^{1}_{\gamma^{\prime\prime}})}q_{s(e^{2}_{\beta^{\prime}})s(f^{2}_{\delta^{\prime\prime}})}\otimes q_{t(f^{1}_{\gamma^{\prime\prime}})t(e^{1}_{\alpha})}q_{s(f^{1}_{\gamma^{\prime\prime}})s(e^{1}_{\alpha})}q_{s(f^{2}_{\delta^{\prime\prime}})s(e^{2}_{\beta^{\prime}})},$
which can be shown to be equal to $\Delta\Big{(}q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}\Big{)}$. $\Box$
Theorem 3.6.
Let $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ be two equivalent triples in the sense of Definition 2.7 so that for $i=1,2$, $\mathcal{G}_{i},\mathcal{H}_{i}$ have the common set of vertices $V$; vertex matrices of $\mathcal{G}_{i}$, $\mathcal{H}_{i}$ are $A_{i},B_{i}$ respectively. Then
${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})\cong{\rm Aut}^{+}(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2}).$
In fact, ${\rm Aut}^{+}(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ can be obtained from ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ by a similarity transformation.
Proof.
Let $T:V\rightarrow V$ be the bijection establishing the equivalence. Then $T_{\ast}:C(V)\rightarrow C(V)$ is an algebra isomorphism. We prove that if $\mathbb{G}$ acts faithfully on the triple $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$, then it also acts faithfully on $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$. As usual we denote the vertex matrices of the graphs $\mathcal{G}_{i}$ and $\mathcal{H}_{i}$ by $A_{i}$, $B_{i}$ respectively; the edge sets of $\mathcal{G}_{i}$ and $\mathcal{H}_{i}$ by $E_{i}$, $F_{i}$ respectively. If $\alpha:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ is the corresponding action then we claim that $\gamma:=(T_{\ast}^{-1}\otimes{\rm id})\circ\alpha\circ T_{\ast}:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ is an action on $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$. Since $T_{\ast}$ is an algebra isomorphism, $\gamma$ is a $C^{\ast}$-action on the $C^{\ast}$-algebra $C(V)$. The corresponding magic unitary for the action $\gamma$ is $T^{-1}UT$ if $U$ is the magic unitary for the action $\alpha$ with respect to some basis of $C(V)$, where $T$ is the permutation matrix with respect to the same choice of basis. As $\mathbb{G}$ is a quantum symmetry of $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$, we have $UA_{1}=A_{1}U$, $UB_{1}=B_{1}U$. As $TA_{2}=A_{1}T$, we have
$(T^{-1}UT)A_{2}=T^{-1}UA_{1}T=A_{2}(T^{-1}UT).$
Similarly using $TB_{2}=B_{1}T$ one can show that $(T^{-1}UT)B_{2}=B_{2}(T^{-1}UT)$. Hence $\gamma$ is a simultaneous quantum symmetry of the graphs $\mathcal{G}_{2},\mathcal{H}_{2}$. So we are left with proving the following commutative diagram:
{tikzcd}
where $\beta$ and $\beta^{\rm op}$ are restrictions of the corepresentation $\gamma^{\otimes 4}$. But it is easy to see that
$\gamma^{\otimes 4}=(T^{-1^{\otimes 4}}_{\ast}\otimes{\rm id})\circ\alpha^{\otimes 4}\circ T_{\ast}^{\otimes 4}.$
Then using the commutation $T^{\otimes 4}_{\ast}\circ\theta_{2\ast}=\theta_{1\ast}\circ T^{\otimes 4}_{\ast}$ and the fact that $\mathbb{G}$ is a quantum symmetry of $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$, we get
$\beta^{\rm op}\circ\theta_{2\ast}$ $=$ $(T^{-1^{\otimes 4}}_{\ast}\otimes{\rm id})\circ\alpha^{\otimes 4}\circ T_{\ast}^{\otimes 4}\circ\theta_{2\ast}$
$=$ $(T^{-1^{\otimes 4}}_{\ast}\otimes{\rm id})\circ\alpha^{\otimes 4}\circ\theta_{1\ast}\circ T_{\ast}^{\otimes 4}$
$=$ $(T^{-1^{\otimes 4}}_{\ast}\otimes{\rm id})\circ(\theta_{1\ast}\otimes{\rm id})\circ\alpha^{\otimes 4}\circ T_{\ast}^{\otimes 4}$
$=$ $(T^{-1^{\otimes 4}}_{\ast}\circ\theta_{1\ast}\otimes{\rm id})\circ\alpha^{\otimes 4}\circ T_{\ast}^{\otimes 4}$
$=$ $(\theta_{2\ast}\otimes{\rm id})\circ(T^{-1^{\otimes 4}}_{\ast}\otimes{\rm id})\circ\alpha^{\otimes 4}\circ T_{\ast}^{\otimes 4}$
$=$ $(\theta_{2\ast}\otimes{\rm id})\circ\beta.$
This proves that $\mathbb{G}$ acts on $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$. Faithfulness is clear. So ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ acts on $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$. Therefore by the universal property, ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ is a quantum subgroup of ${\rm Aut}^{+}(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$. Similarly considering $T^{-1}$, one can show that ${\rm Aut}^{+}(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ is a quantum subgroup of ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ i.e.
${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})\cong{\rm Aut}^{+}(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2}).$
The assertion about the similarity transformation is clear. $\Box$
Now we are ready to define the notion of the quantum automorphism group of a $2$-graph.
Definition 3.7.
Let $\Lambda$ be a $2$-graph with the defining triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$. Then the quantum automorphism group of $\Lambda$ is defined to be the quantum automorphism group of $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ as defined in Definition 3.4.
Remark 3.8.
As explained earlier, thanks to Theorem 3.6 and Theorem A.1, the notion of the quantum automorphism group of a $2$-graph $\Lambda$ with the defining triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ is well-defined. It will be denoted by either of the notations ${\rm Aut}^{+}(\Lambda)$ and ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$.
4. Computations
(a) Two copies of the same graph with the trivial bijection:
Recall the pullback of a $1$-graph $\mathcal{G}$ via the monoid morphism $f:\mathbb{N}^{2}\rightarrow\mathbb{N}$ given by $f(m,n)=m+n$. The pullback $f^{\ast}(\mathcal{G})$ is the $2$-graph with the defining triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$, where $\mathcal{G}_{i}$’s are two copies of the same graph $\mathcal{G}$ and the bijection $\theta$ between the composable pairs is the trivial bijection. Since $\theta$ is the trivial bijection, replacing $(e^{2}_{\mu},e^{1}_{\nu})$ by $(e^{1}_{\alpha},e^{2}_{\beta})$ and $(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$ by $(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})$ in the Equations (3.3), we see that (3.3) are redundant. Also since the individual $1$-graphs are two copies of the same graph, the following is immediate:
${\rm Aut}^{+}(f^{\ast}(\mathcal{G}))\cong{\rm Aut}^{+}(\mathcal{G}).$
Thus we see that the quantum automorphism group of any $1$-graph can be realized as the quantum automorphism group of a $2$-graph $\Lambda$. Consequently we always have examples of $2$-graphs admitting genuine quantum automorphism groups.
(b) Two different graphs with a nontrivial bijection:
We take the following $1$-graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ as the constituent graphs to construct the $2$-graph whose $1$-skeleton is given as follows:
Graph $\mathcal{G}_{1}$ Graph $\mathcal{G}_{2}$ 1-Skeleton
Denote the edge sets of $\mathcal{G}_{1},\mathcal{G}_{2}$ by $E_{1},E_{2}$ respectively. Then the composable pairs are given by
$E_{1}\star E_{2}=\{(e_{1},f_{1}),(e_{2},f_{3}),(e_{3},f_{2}),(e_{4},f_{4})\},$
$E_{2}\star E_{1}=\{(f_{1},e_{4}),(f_{2},e_{2}),(f_{3},e_{3}),(f_{4},e_{1})\}.$
We take the following choice of $\theta$:
$(e_{1},f_{1})\mapsto(f_{3},e_{3})$
$(e_{2},f_{3})\mapsto(f_{1},e_{4})$
$(e_{3},f_{2})\mapsto(f_{4},e_{1})$
$(e_{4},f_{4})\mapsto(f_{2},e_{2})$
In fact the reader can verify that in this example there is a unique choice of $\theta$ given as above. The vertex matrices of the above two graphs $\mathcal{G}_{1},\mathcal{G}_{2}$ are respectively
$A_{1}=\begin{pmatrix}0&1&0&0\\ 1&0&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}$; $A_{2}=\begin{pmatrix}0&0&0&1\\ 0&0&1&0\\ 0&1&0&0\\ 1&0&0&0\end{pmatrix}$.
With the usual notations, the action $\alpha$ of ${\rm Aut}^{+}(\mathcal{G}_{1}.\mathcal{G}_{2},\theta)$ on $C(V)$ is given by $\alpha(\delta_{i})=\sum_{j=1}^{4}\delta_{j}\otimes q_{ji}$. Then the matrix $((q_{ij}))$ commutes with both the vertex matrices $A_{1},A_{2}$. A simple calculation reveals that the commutations impose the following relations among the generators $q_{ij}$:
(4.1) $q_{11}=q_{22}=q_{33}=q_{44};\ q_{12}=q_{21}=q_{34}=q_{43};$
(4.2) $q_{14}=q_{23}=q_{32}=q_{41};\ q_{13}=q_{31}=q_{24}=q_{42}.$
We claim that the relations (3.3) are redundant for the graph $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$. We shall prove a somewhat stronger result: For any choice of edges from $\mathcal{G}_{2}$, the first two terms of the LHS of (3.3) are equal i.e. $q_{t(f_{i})t(f_{j})}q_{s(f_{i})s(f_{j})}=q_{t(f_{i})t(f_{j})}$ for all $i,j=1,2,3,4$. This can be checked by direct calculation using the relations (4.1) and (4.2). For example:
$q_{t(f_{1})t(f_{2})}q_{s(f_{1})s(f_{2})}=q_{14}q_{41}=q_{14};$
$q_{t(f_{1})t(f_{1})}q_{s(f_{1})s(f_{1})}=q_{11}q_{44}=q_{44}.$
Similarly, one can prove that for $i,j=1,2,3,4$, $q_{t(e_{i})t(e_{j})}q_{s(e_{i})s(e_{j})}=q_{t(e_{i})t(e_{j})}$ so that, in particular, for all $(e^{1}_{\alpha},e^{2}_{\beta})\in E_{1}\star E_{2}$ and all $(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})\in E_{2}\star E_{1}$ with $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$ and $\theta(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})=(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$,
$q_{t(e^{2}_{\mu^{\prime}})t(e^{2}_{\mu})}q_{s(e^{2}_{\mu^{\prime}})s(e^{2}_{\mu})}q_{s(e^{1}_{\nu^{\prime}})s(e^{1}_{\nu})}$
$=$ $q_{t(e^{2}_{\mu^{\prime}})t(e^{2}_{\mu})}q_{s(e^{1}_{\nu^{\prime}})s(e^{1}_{\nu})}$
$=$ $q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}$
$=$ $q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}$
Hence the fundamental unitary matrix of ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ reduces to
$\begin{pmatrix}q_{1}\ q_{2}\ q_{3}\ q_{4}\\ q_{2}\ q_{1}\ q_{4}\ q_{3}\\ q_{3}\ q_{4}\ q_{1}\ q_{2}\\ q_{4}\ q_{3}\ q_{2}\ q_{1}\\ \end{pmatrix}$,
where $q_{i}$’s are mutually orthogonal projections such that $\sum_{i=1}^{4}q_{i}=1$. Now it is easy to check that the above fundamental unitary corresponds to the CQG $C^{\ast}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$. Therefore ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ is isomorphic to the classical group $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$.
(c) Two copies of the graph $\mathcal{G}$ where $\mathcal{G}$ is the complete graph on four vertices:
{tikzpicture}
We denote one set of edges by $e_{i}$’s and the other set by $f_{i}$’s. The composable pairs are given by
$E\ \star\ E=\left\{\begin{matrix}&(e_{1},f_{2}),&(e_{1},f_{10}),&(e_{1},f_{7}),&(e_{2},f_{4}),&(e_{2},f_{12}),&(e_{2},f_{1}),\\ &(e_{3},f_{4}),&(e_{3},f_{12}),&(e_{3},f_{1}),&(e_{4},f_{3}),&(e_{4},f_{9}),&(e_{4},f_{5}),\\ &(e_{5},f_{6}),&(e_{5},f_{11}),&(e_{5},f_{8}),&(e_{6},f_{5}),&(e_{6},f_{3}),&(e_{6},f_{9}),\\ &(e_{7},f_{6}),&(e_{7},f_{11}),&(e_{7},f_{8}),&(e_{8},f_{2}),&(e_{8},f_{1}0),&(e_{8},f_{7}),\\ &(e_{9},f_{2}),&(e_{9},f_{10}),&(e_{9},f_{7}),&(e_{10},f_{3}),&(e_{10},f_{9}),&(e_{10},f5),\\ &(e_{11},f_{1}),&(e_{11},f_{4}),&(e_{11},f_{12}),&(e_{12},f_{6}),&(e_{12},f_{11}),&(e_{12},f_{8})\end{matrix}\right\}.$
We choose the map $\theta$ to be the following:
$(e_{1},f_{2})\mapsto(e_{4},f_{3}),$
$(e_{4},f_{3})\mapsto(e_{1},f_{2}),$
$(e_{8},f_{7})\mapsto(e_{6},f_{5}),$
$(e_{6},f_{5})\mapsto(e_{8},f_{7}),$
keeping rest of the elements of $E\star E$ fixed. As usual we write the action of ${\rm Aut}^{+}(\mathcal{G},\mathcal{G},\theta)$ on the set of vertices by
$\alpha(\delta_{i})=\sum_{j=1}^{4}\delta_{j}\otimes q_{ji},\ {\rm for}\ i=1,\ldots,4.$
We assert that
  1. (1)
    $q_{13}=q_{12}=0$ (implies $q_{31}=q_{21}=0$),
  2. (2)
    $q_{43}=q_{42}=0$ (implies $q_{34}=q_{24}=0$).
To prove our assertions, we shall first establish the equality $q_{13}=q_{12}$. By subsequently multiplying both sides of this equation by $q_{13}$, we deduce that $q_{13}$ equals zero. To accomplish this, we show that $q_{i1}q_{13}q_{j1}=q_{i1}q_{12}q_{j1}$ for all $i,j=1,2,\ldots,4$. Recalling from equation (3.3) for $\theta(e,f)=(f^{\prime},e^{\prime})$ and $\theta(\mu,\nu)=(\nu^{\prime},\mu^{\prime})$ we have
$q_{t(\mu)t(e)}q_{s(\mu)s(e)}q_{s(\nu)s(f)}=q_{t(\nu^{\prime})t(f^{\prime})}q_{s(\nu^{\prime})s(f^{\prime})}q_{s(\mu^{\prime})s(e^{\prime})}.$
For $(e,f)=(e_{1},f_{2})$ and $(f^{\prime},e^{\prime})=(e_{4},f_{3})$ the above equation reduces to
$q_{t(\mu)1}q_{s(\mu)3}q_{s(\nu)1}=q_{t(\nu^{\prime})1}q_{s(\nu^{\prime})2}q_{t(\mu^{\prime})1}.$
We have the following equations for different choices of $(\mu,\nu)\in\{(e_{3},f_{4}),(e_{3},f_{1}),\\ (e_{3},f_{12}),(e_{2},f_{4}),(e_{2},f_{1}),(e_{2},f_{12}),(e_{11},f_{4}),(e_{11},f_{1}),(e_{11},f_{12})\}$:
$q_{21}q_{13}q_{21}=q_{21}q_{12}q_{21}$, $q_{21}q_{13}q_{31}=q_{21}q_{12}q_{31}$, $q_{21}q_{13}q_{41}=q_{21}q_{12}q_{41}$, $q_{31}q_{13}q_{21}=q_{31}q_{12}q_{21}$, $q_{31}q_{13}q_{31}=q_{31}q_{12}q_{31}$, $q_{31}q_{13}q_{41}=q_{31}q_{12}q_{41}$, $q_{41}q_{13}q_{21}=q_{41}q_{12}q_{21}$, $q_{41}q_{13}q_{31}=q_{41}q_{12}q_{31}$, $q_{41}q_{13}q_{41}=q_{41}q_{12}q_{41}$,
whereas the other equations follow automatically as both sides of the equations are 00 using the orthogonality properties of $q_{ij}$’s. Adding up all the equations above we get
$\sum_{i,j=1}^{4}q_{i1}q_{13}q_{j1}=\sum_{i,j=1}^{4}q_{i1}q_{12}q_{j1}$
which clearly shows that $q_{13}=q_{12}$. A similar calculation shows that $q_{43}=q_{42}$. Therefore the fundamental matrix takes the following form:
$\left(\begin{matrix}q_{11}&0&0&q_{14}\\ 0&q_{22}&q_{23}&0\\ 0&q_{32}&q_{33}&0\\ q_{41}&0&0&q_{44}\end{matrix}\right).$
Assume that $q_{11}=p$ and $q_{22}=q$, where $p$ and $q$ are projections. Then the magic condition forces $q_{41}=1-p=q_{14},q_{44}=p,q_{32}=1-q=q_{23},q_{33}=q$.
In the above calculation putting $(\mu,\nu)=(e_{8},f_{2})=(\nu^{\prime},\mu^{\prime})$ we see that
$q_{41}q_{33}q_{11}=q_{41}q_{32}q_{11}.$
Hence we have $(1-p)qp=(1-p)(1-q)p$ which implies $qp=pqp$. Consequently, $qp=pqp=(qp)^{*}=pq$ shows the $C^{*}$-algebra generated by $q_{ij}$’s is commutative. Therefore, the quantum automorphism group is actually classical and in fact a subgroup of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. Now it can be checked easily that none of the order $2$ elements in $S_{4}$ generating $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ belongs to the symmetry group. Hence the quantum automorphism group is not only classical, but also trivial.
Concluding remarks
  • The $2$-graph in the example (b) can be seen to be isomorphic to the cartesian product of the $1$-graph with two vertices and a bidirected edge (say $\mathcal{G}$) with itself in the sense of Proposition 1.8 of [NYJM]. In that example, the quantum automorphism group of the $2$-graph is ${\rm Aut}^{+}(\mathcal{G})\otimes{\rm Aut}^{+}(\mathcal{G})$. We conjecture that if $\mathcal{G}_{1},\mathcal{G}_{2}$ are two $1$-graphs, then the quantum automorphism group of the $2$-graph $\mathcal{G}_{1}\times\mathcal{G}_{2}$ at least contains ${\rm Aut}^{+}(\mathcal{G}_{1})\otimes{\rm Aut}^{+}(\mathcal{G}_{2})$ as a quantum subgroup.
  • We believe that the idea in this paper can be adapted to define the notion of the quantum automorphism group of $k$-graphs in general as a $k$-graph can be described by similar combinatorial data of $k$-colored graphs (see [Hazzlewood1]). But we believe that the computations will get harder as $k$ increases.
  • The quantum automorphism group of a $2$-graph should naturally act on the corresponding higher rank graph $C^{\ast}$-algebra as it has already been observed in the literature in the case of $1$-graphs (see [Joardar1],[Joardar2],[Weber] for example).
  • The quantum automorphism group of a $2$-graph has natural corepresentations on the morphism spaces of the $2$-graph. Therefore our consideration might also shed some light on the corepresentation theory of the quantum automorphism groups of $2$-graphs.
Appendix A
Theorem A.1.
Two defining triples $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ are equivalent if and only if the corresponding 2-graphs (say $\Lambda_{1}$ and $\Lambda_{2}$ respectively) are isomorphic.
Proof.
This theorem must be well-known and proof being very natural, is only sketched here. Suppose $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ are equivalent. Therefore both $(\mathcal{G}_{i},\mathcal{H}_{i},\theta_{i})$ for $i=1,2$ share the common vertex set $V$; a bijection $T:V\rightarrow V$; $\mathcal{G}_{i},\mathcal{H}_{i}$ have adjacency matrices $A_{i},B_{i}$ respectively for $i=1,2$ such that
$TA_{2}=A_{1}T;\ TB_{2}=B_{1}T.$
Let’s construct a functor $\Phi$ from $\Lambda_{1}$ to $\Lambda_{2}$. We recall that ${\rm obj}(\Lambda_{i})$ are identified with the common vertex set $V$. We assign to each $i\in V$ the element $T(i)\in V$. So that we associate each object of $\Lambda_{1}$ to an object in $\Lambda_{2}$. Now we want to associate each morphism in $\Lambda_{1}^{(m,n)}$ to a morphism in $\Lambda_{2}^{(m,n)}$ for all $(m,n)\in\mathbb{N}^{2}$. For $(m,n)=(1,0)$, a morphism $f$ from an object $i$ to an object $j$ in $\Lambda_{1}$ is an edge (say $f$) in the graph $\mathcal{G}_{1}$ with $s(f)=i$ and $r(f)=j$. Since $TA_{2}=A_{1}T$, there is an edge say $\Phi(f)$ with $s(\Phi(f))=T(i)$ and $r(\Phi(f))=T(j)$. Hence we get the following commutative diagram where $f\in\Lambda_{1}^{(1,0)}$ and $\Phi(f)\in\Lambda_{2}^{(1,0)}$:
{tikzcd}
Similarly one gets a commutative diagram involving morphisms of degree $(0,1)$. Now let $F$ be a morphism of degree $(1,1)$ with $s(F)=i$ and $r(F)=j$. By the unique factorization property, there are unique morphisms $f_{1},f_{2}$ of degree $(1,0)$ and $g_{1},g_{2}$ of degree $(0,1)$ such that $F=f_{1}\circ g_{1}=g_{2}\circ f_{2}$. It is clear that two morphisms $f,g$ of degree $(1,0)$ or $(0,1)$ are composable iff $\Phi(f)$ and $\Phi(g)$ are composable. Then we define $\Phi(F)$ to be either of $\Phi(f_{1})\circ\Phi(g_{1})$ or $\Phi(g_{2})\circ\Phi(f_{2})$. Using the $\theta$-compatibility and the fact that $\theta_{1}(f_{1},g_{1})=(g_{2},f_{2})$ we get that $\theta_{2}(\Phi(f_{1}),\Phi(g_{1}))=(\Phi(g_{2}),\Phi(f_{2}))$. This proves the well-definedness of $\Phi(F)$. Now step by step one can define $\Phi(F)\in\Lambda_{2}^{(m,n)}$ for any $F\in\Lambda_{1}^{(m,n)}$ and for any $(m,n)\in\mathbb{N}^{2}$ and one easily verifies that $\Phi$ is a functor. Similarly one can define a functor $\Phi^{-1}$ from $\Lambda_{2}$ to $\Lambda_{1}$ using the map $T^{-1}$ on the vertex set $V$ (i.e. on the level of objects) and it is easily verified that $\Phi^{-1}$ is the inverse of the functor $\Phi$. Thus an isomorphism between the categories $\Lambda_{1}$ and $\Lambda_{2}$ is established.
For the converse direction, we start with two isomorphic $2$-graphs $\Lambda_{1}$ and $\Lambda_{2}$ with defining triples $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ respectively. Using the isomorphism of $\Lambda_{1},\Lambda_{2}$, one can see that there is a bijection between ${\rm obj}(\Lambda_{1})$ and ${\rm obj}(\Lambda_{2})$ which are finite sets by our assumption. So one can without loss of generality assume that $\mathcal{G}_{i},\mathcal{H}_{i}$ share the same vertex set $V$. We denote the functor from $\Lambda_{1}$ to $\Lambda_{2}$ by $\Phi$. Then there is a morphism of degree $(1,0)$ from object $v_{i}$ to $v_{j}$ i.e. there is an edge of the graph $\mathcal{G}_{1}$ say $f$ from the vertex $v_{i}$ to the vertex $v_{j}$Rightarrow there is a unique morphism $\Phi(f)$ from $\Phi(v_{i})$ to $\Phi(v_{j})$ i.e. there is an edge from the vertex $\Phi(v_{i})$ to the vertex $\Phi(v_{j})$. Equivalently if the permutation matrix corresponding to the bijection $\Phi$ is denoted by $T$, $TA_{2}=A_{1}T$ where $A_{i}$ are the vertex matrices of $\mathcal{G}_{i}$ for $i=1,2$. Similarly considering degree $(0,1)$ morphisms one gets $TB_{2}=B_{1}T$. Let $F$ be a morphism of degree $(1,1)$ in $\Lambda_{1}$ with unique factorization $f_{1}\circ g_{1}=g_{1}^{\prime}\circ f_{1}^{\prime}$. Then, as usual, identifying morphisms of degree $(1,0)$ and $(0,1)$ in $\Lambda_{1}$ with edges of graphs $\mathcal{G}_{1},\mathcal{H}_{1}$ respectively, we get $\theta_{1}(f_{1},g_{1})=(g_{1}^{\prime},f_{1}^{\prime})$. But by the definition of functor, $\Phi(F)=\Phi(f_{1})\circ\Phi(g_{1})=\Phi(g_{1}^{\prime})\circ\Phi(f_{1}^{\prime})$. Therefore from the unique factorization of $\Phi(F)$ in $\Lambda_{2}$, we obtain that $\theta_{2}(\Phi(f_{1}),\Phi(g_{1}))=(\Phi(g_{1}^{\prime}),\Phi(f_{1}^{\prime}))$ so that we get the required $\theta$ compatibility to ensure that the triples $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ are equivalent. $\Box$
Acknowledgement The first author is supported by SERB MATRICS grant no. MTR/2022/000515 (Govt. of India). The second author is partially supported by JC Bose Fellowship and grant of Prof Debashish Goswami, ISI Kolkata, awarded to him by SERB, DST ( Govt of India ). Both the authors would like to thank Dr. Arnab Mandal for some useful discussions.